# Evaluating the Capability of ChatGPT on Ancient Chinese

### 23 Dec 2023 v1

@@ Abstract
ChatGPT’s proficiency in handling modern standard languages suggests potential for its use in understanding ancient Chinese. This project explores ChatGPT’s capabilities on ancient Chinese via two tasks: translating ancient Chinese to modern Chinese and recognizing ancient Chinese names. A comparison of ChatGPT’s output with human translations serves to evaluate its comprehension of ancient Chinese. The findings indicate that: (1.)the proficiency of ancient Chinese by ChatGPT is yet to reach a satisfactory level; (2.) ChatGPT performs the best on ancient-to-modern translation when feeding with three context sentences. To help reproduce our work, we display the python code snippets used in this study.
**footnotetext: The corresponding author is Shijing Si, shijing.si@outlook.com.
1 Introduction
@@
Ancient languages serve as repositories for humanity’s historical and cultural heritage. As a treasure and important heritage of ancient Chinese culture (Lee and Wong,, 2012), ancient Chinese has important cultural significance and historical value in the study of natural language processing.
@@
Recently there have been considerable advancements made in the analysis and interpretation of ancient Chinese through the use of deep learning models (Cheng et al.,, 2020; Tian et al.,, 2021; Tang et al.,, 2022; Anderson et al.,, 2023; Wang et al., 2023c, ). However, the understanding of ancient Chinese is a very challenging task in natural language processing due to its complex grammatical structures, cultural nuances, and polysemy of the language (Yu and Huangfu,, 2019; Jin et al.,, 2023; Guo et al.,, 2023).
@@
Deep learning methods has achieved significant progress in many natural language processing (NLP) tasks (Si et al.,, 2019; Devlin et al.,, 2018; Si et al.,, 2020; Brown et al.,, 2020). Notably, large language models like ChatGPT have been shown powerful generation and understanding capabilities across many languages (Sundararaman et al.,, 2020; Lai et al.,, 2023; Liu et al.,, 2023; Ray,, 2023; Bian et al.,, 2023), and has attracted wide attention and application around the world since its release (Fang et al.,, 2023; Wang and Demszky,, 2023; Wu et al., 2023b, ; Kocoń et al.,, 2023; Gilardi et al.,, 2023; Ronanki et al.,, 2023).
@@
However, there are few research on the use of ChatGPT for ancient Chinese processing. Jin et al., (2023) evaluates the performance of a few machine translation methods including ChatGPT. Lan, (2023) systematically examined ChatGPT at six cognitive levels on the Book of changes (易经). Following these works, we study the capability of ChatGPT on ancient-to-modern Chinese translation and people name recognition (Zhang et al.,, 2021) in ancient Chinese text. We leveraged the ancient Chinese book, called Shi Shuo Xin Yu(世说新语), to assess ChatGPT’s ability in understanding ancient Chinese through two tasks: the task of translating ancient Chinese into modern Chinese and the task of recognizing people’s names in ancient Chinese.
@@
In this paper, we evaluate the capability of ChatGPT on an ancient Chinese book, Shi Shuo Xin Yu, which is largely ignored by previous research. Also we study the performance of ChatGPT on ancient-to-modern translation by varying the input length for each individual query. Additionally, the personal name recognition is rarely explored in the ancient Chinese processing.
@@
The article is organized as follows. Section 2 presents the related works on ChatGPT and ancient Chinese processing. Section 3 demonstrates the experimental setup and implementation code. Section 4 exhibits the experimental results and analysis. Section 5 shows the conclusion and discussion.
2 Related Works
@@
ChatGPT and other large language models have indeed garnered significant attention due to their impressive capabilities. However, these large model are not perfect and still have limitations, including producing inconsistent or nonsensical responses, or writing information that seems factual but is actually completely made up. They are sensitive to the way prompts are phrased, and don’t possess any real understanding of the world or self-awareness. Therefore, there are many studies that explore and evaluate the performance of ChatGPT on many tasks and examinations. We will illustrate related works in three subsections.
2.1 Application of ChatGPT in Natural Language Processing
@@
Ogundare and Araya, (2023) utilized spontaneous quality (SQ) scores to compare the performance of ChatGPT on many NLP tasks such as machine translation, machine summarization, question answering, and language generation, compared with other mainstream algorithms. Wang et al., 2023b , taking document-level machine translation (MT) as an experimental platform, conducts an in-depth evaluation of ChatGPT’s discourse modeling ability in three aspects: the role of discourse awareness cue, the comparison of translation models, and the analysis of discourse modeling ability. Zhang et al., (2023) explores ChatGPT’s ability to summarize and evaluate on cross-lingual cross-time summarization (CLCTS) tasks. Ronanki et al., (2023) explores the application of ChatGPT for user story quality assessment and compares its performance to existing benchmarks. Sun et al., (2023) proposed a set of generic modules, attempting to break through the limitations of ChatGPT on various NLP tasks, including question answering, common sense reasoning, natural language reasoning, etc. Antoun et al., (2023) proposed a method for developing and evaluating ChatGPT detectors for French text, focusing on their robustness against extra-territorially based data and common attack schemes. Wang and Demszky, (2023) explore whether ChatGPT can be a cost-effective supplement to expert feedback by acting as an automated teacher or coach.
2.2 Application of ChatGPT on Other Fields
@@
Many research explore the ability of ChatGPT to understand texts in other fields than NLP. Cao et al., (2023) investigated ChatGPT’s potential cultural background by analyzing its answers to questions aimed at quantifying human cultural differences. The results of the study showed that when inspired by an American context, ChatGPT showed strong consistency with American culture, but it was less effective in adapting to other cultural contexts. George and George, (2023) explore how ChatGPT can enhance e-commerce through chat and other sectors such as education, entertainment, finance, health, news, and productivity, and how ChatGPT can be used to create more personalized content for users and help businesses improve the efficiency and effectiveness of customer service. Rezayi et al., (2023) explored a new field of agricultural natural language processing by studying the effectiveness of pre-training transformer-based language models using food-related text corpora. Wang et al., 2023d investigated ChatGPT’s ability to infer dynamic network structures from temporal text data, especially financial news. Khondaker et al., (2023) conducted a large-scale evaluation of ChatGPT on a wide range of Arabic NLP tasks. Wu et al., 2023a evaluated ChatGPT’s ability in knowledge in the medical field on the Chinese National Medical Licensing Examination (CNMLE).
2.3 Application of Large Language Models in Ancient Chinese
@@
However, there is relatively little research on ChatGPT’s understanding of ancient Chinese. The research on the understanding of ancient Chinese by language model is mainly based on BERT model. Wang et al., (2022) and Chang et al., (2023) built a pre-trained language model of SikuBERT and SikuGPT for ancient text intelligent processing tasks based on BERT deep language model framework using the verified high-quality full-text corpus of SikukuQuanshu as a training set. Some researchers built GuwenBERT based on RoBERTa model trained on a large number of ancient literature corpus. Tsinghua University has constructed BERT CCPoem, a pre-training model based on BERT specifically for ancient Chinese poetry, which is trained on the complete collection of ancient Chinese poetry CCPC full v1.0 and can be used for intelligent poetry retrieval, recommendation and sentiment analysis. Wang et al., 2023a introduced two language models, GujiBERT and GujiGPT, which are foundational models specifically designed for intelligent information processing of ancient texts.
3 Experiments
@@
3.1 Data preparation
The main data comes from Github open source project NiuTrans/classic-Modern***The resource can be found in website:https://github.com/NiuTrans/Classical-Modern, which is a very comprehensive ancient Chinese-modern parallel corpus, including a large number of classical ancient works. This dataset divides and displays each ancient book by chapters, and the text is stored in the txt documents under each chapter. These bilingual data are provided in sentence level, and three data formats are provided: original text, translation and bilingual.
In this study, the team mainly utilized the bilingual data of the book, Shi Shuo Xin Yu (世说新语), including 36 chapters, to satisfy our needs in the experiments. And most articles in this book are about biographies of famous people in ancient China, which can be used for ancient-to-modern Chinese translation and people’s name recognition.
Due to the input length limitation of ChatGPT, we split the whole book into 3923 sentences and each time we feed the model a sentence. For the ancient-to-modern Chinese translation task, we include all sentences, but for the ancient Chinese name recognition task, we randomly select a total of 300 sentences because of time constraint. Also we manually label Chinese names for the 300 sentences. The data is shown in Table 1 as fowllows.
Table 1: A few examples used in the experiments. The sentences are extracted from Shi Shuo Xin Yu, and Chinese names are manually labeled by the authors.
@@ 3.2 Accesss to ChatGPT
3.3 Translation of ancient Chinese by ChatGPT
Once having access to ChatGPT in Python, we can simply design a sensible prompt to perform the ancient-to-modern Chinese translation task. We tried a variety of prompt input methods, such as ”Please translate the following ancient Chinese:”, ”Translate the ancient Chinese after the colon into modern Chinese:”, ”Translate the ancient Chinese into modern Chinese” and so on. We found that it is necessary to include ”translation” and ”modern Chinese” in order to perform the ancient-to-modern Chinese translation robustly, otherwise the output of ChatGPT may have redundant and wrong information. In addition, we also found that even if we input only ancient Chinese sentences without any instruction, ChatGPT can sometimes return the correct translation results in modern Chinese.
After many trials, we choose a straightforward ”将文言文翻译成现代汉语(translate ancient Chinese into modern Chinese)” as the prompt instruction for the translation task. This prompt almost never produces error information, and no superfluous information. we extract the text content that contains only the modern Chinese translation of the ancient Chinese. The code snippet used for the ancient-to-modern Chinese translation is displayed in Fig. 1.
@@ 3.3 Translation of ancient Chinese by ChatGPT
Due to the input length constraints set by OpenAI, it is unfeasible to command ChatGPT to translate the entire Shi Shuo Xin Yu book directly. Consequently, we use a systematic FOR loop to sequentially translate preloaded segments of the book. To proportionately evaluate its performance, we adjust the size of the input - the amount of text that needs to be translated. We start off with a single sentence to observe how well the program can perform a basic translation task. Then, they slowly increase the input length to three, five, and eight sentences. This gradual scaling may challenge the program and test its limits, especially in maintaining cohesiveness and continuity in the translation of longer passages.
For the ancient-to-modern translation task, we evaluate the performance of ChatGPT with BLEU (Papineni et al.,, 2002) and BERT-Score(Zhang et al.,, 2019).
3.4 People Name Recognition from Ancient Chinese Literature
Besides translation, we assess the competence of ChatGPT in recognizing names referenced within the ancient Chinese text. It is noteworthy that the effectiveness of ChatGPT in executing this task is contingent upon the prompt instruction. We utilize a trial-and-error approach to ascertain a suitable prompt instruction, ”已知文言文的原文如下,找出该文言文中的人名”, which means that ”Given the following ancient Chinese text, find the people names in it”. The code snippet for people name recognition from ancient Chinese text is exhibited in Fig. 2.
@@ 3.4 People Name Recognition from Ancient Chinese Literature
The Jieba module in Python is widely employed for word segmentation and rudimentary named entity recognition in Chinese texts. Hence, for our study on the recognition of people names in ancient Chinese text, we incorporate Jieba as a baseline method for comparison. For the people name recognition task, we evaluate the performance of ChatGPT with commonly used metrics, like precision, recall and F1 score (Jiang et al.,, 2016).
4 Results and Analysis
@@
4.1 Assessing the Efficacy of ChatGPT in the Translation of Ancient Chinese Texts
As mentioned in the experimental setup section 3, we deploy ChatGPT to convert sentences from the ancient Chinese text, Shi Shuo Xin Yu, into modern Chinese in configurations of 1, 3, 5, and 8 sentences per query. These translations are then evaluated and contrasted using quantitative metrics such as BLEU, and BERT-Score. Table 2 illustrates the BLEU scores resulting from the translation of ancient to modern Chinese from ChatGPT under four different configurations: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per individual query. In this table, we conduct an evaluation of the BLEU scores across three different levels of granularity: 1-gram, 2-gram, and 4-gram. From Table 2, it is observed that ChatGPT renders the most proficient translations from ancient to modern Chinese when it is engaged with three sequential ancient Chinese sentences (3-sent.), garnering 22.60% at 2-gram and 20.19‰ at 4-gram Bleu scores. This suggests that a shorter contextual input enhances the translation quality from ancient to modern Chinese. Interestingly, the quality of translation does not exhibit improvement when the context increases to eight sentences, mirroring the performance with a single sentence input. Another important finding is that the BLUE scores are still very low, reaching 0 at 4-gram BLUE in all four settings, which means that ChatGPT with GPT-3.5-Turbo still performs very poorly in ancient-to-modern Chinese translation.
Table 2: The 1-gram, 2-gram, and 4-gram BLUE scores of translated modern Chinese texts from ChatGPT under four settings: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per query.

### 23 Feb 2024 v2

# Exploring the Capabilities of ChatGPT in Ancient Chinese Translation and Person Name Recognition

@@ Abstract
ChatGPT’s proficiency in handling modern standard languages suggests potential for its use in understanding ancient Chinese. This paper explores ChatGPT’s capabilities on ancient Chinese via two tasks: translating ancient Chinese to modern Chinese and recognizing ancient Chinese names. A comparison of ChatGPT’s output with human translations serves to evaluate its comprehension of ancient Chinese. The findings indicate that: (1.)the proficiency of ancient Chinese by ChatGPT is yet to reach a satisfactory level; (2.) ChatGPT performs the best on ancient-to-modern translation when feeding with three context sentences. To help reproduce our work, we display the python code snippets used in this study.
Index Terms:
Large Language Model, ChatGPT, Machine Translation, Ancient Chinese, Person Name Recognition
I Introduction
@@
Ancient languages serve as repositories for humanity’s historical and cultural heritage. As a treasure and important heritage of ancient Chinese culture (Lee and Wong, 2012), ancient Chinese has important cultural significance and historical value in the study of natural language processing.
@@
Recently there have been considerable advancements made in the analysis and interpretation of ancient Chinese through the use of deep learning models (Cheng et al., 2020; Tian et al., 2021; Tang et al., 2022; Anderson et al., 2023; Wang et al., 2023a). However, the understanding of ancient Chinese is a very challenging task in natural language processing due to its complex grammatical structures, cultural nuances, and polysemy of the language (Yu and Huangfu, 2019; Jin et al., 2023; Guo et al., 2023).
@@
Deep learning methods has achieved significant progress in many natural language processing (NLP) tasks (Si et al., 2019; Devlin et al., 2018; Si et al., 2020; Brown et al., 2020). Notably, large language models like ChatGPT have been shown powerful generation and understanding capabilities across many languages (Sundararaman et al., 2020; Lai et al., 2023; Liu et al., 2023; Ray, 2023; Bian et al., 2023), and has attracted wide attention and application around the world since its release (Fang et al., 2023; Wang and Demszky, 2023; Wu et al., 2023a; Kocoń et al., 2023; Gilardi et al., 2023; Ronanki et al., 2023).
@@
However, there are few research on the use of ChatGPT for ancient Chinese processing. Jin et al. (2023) evaluates the performance of a few machine translation methods including ChatGPT. Lan (2023) systematically examined ChatGPT at six cognitive levels on the Book of changes (易经, Yi Jing in Chinese). Following these works, we study the capability of ChatGPT on ancient-to-modern Chinese translation and people name recognition (Zhang et al., 2021) in ancient Chinese text. We leveraged the ancient Chinese book, called Shi Shuo Xin Yu(世说新语 in Chinese characters), to assess ChatGPT’s ability in understanding ancient Chinese through two tasks: the task of translating ancient Chinese into modern Chinese and the task of recognizing people’s names in ancient Chinese.
@@
In this paper, we evaluate the capability of ChatGPT on an ancient Chinese book, Shi Shuo Xin Yu, which is largely ignored by previous research. Also we study the performance of ChatGPT on ancient-to-modern translation by varying the input length for each individual query. Additionally, the personal name recognition is rarely explored in the ancient Chinese processing. To summarize, our cntributions are shown as follows:
  • We evaluate the ancient-to-modern Chinese translation of ChatGPT on Shi Shuo Xin Yu, and observe that the translation quality depends on the context length. From our experiments, ChatGPT performs the best on ancient-to-modern translation when three consecutive sentences are fed into the prompt instruction.
  • We also investigate the capability of ChatGPT on ancient person name recognition, finding that ChatGPT outperforms the commonly used method in Jieba, and its performance improves as the number of demonstrations increases.
@@
The article is organized as follows. Section II presents the related works on ChatGPT and ancient Chinese processing. Section III demonstrates the experimental setup and implementation code. Section IV exhibits the experimental results and analysis. Section V shows the conclusion and discussion.
II Related Works
@@
ChatGPT and other large language models have indeed garnered significant attention due to their impressive capabilities Ciesla (2024). However, these large model are not perfect and still have limitations, including producing inconsistent or nonsensical responses, or writing information that seems factual but is actually completely made up. They are sensitive to the way prompts are phrased, and fail to possess any real understanding of the world or self-awareness. Therefore, there are many studies that explore and evaluate the performance of ChatGPT on many tasks and examinations. We will illustrate related works in three subsections.
II-A Application of ChatGPT in Natural Language Processing
@@
Ogundare and Araya (2023) utilized spontaneous quality (SQ) scores to compare the performance of ChatGPT on many NLP tasks such as machine translation, machine summarization, question answering, and language generation, compared with other mainstream algorithms. Wang et al. (2023b), taking document-level machine translation (MT) as an experimental platform, conducts an in-depth evaluation of ChatGPT’s discourse modeling ability in three aspects: the role of discourse awareness cue, the comparison of translation models, and the analysis of discourse modeling ability. Zhang et al. (2023) explores ChatGPT’s ability to summarize and evaluate on cross-lingual cross-time summarization (CLCTS) tasks. Ronanki et al. (2023) explores the application of ChatGPT for user story quality assessment and compares its performance to existing benchmarks. Sun et al. (2023) proposed a set of generic modules, attempting to break through the limitations of ChatGPT on various NLP tasks, including question answering, common sense reasoning, natural language reasoning, etc. Antoun et al. (2023) proposed a method for developing and evaluating ChatGPT detectors for French text, focusing on their robustness against extra-territorially based data and common attack schemes. Wang and Demszky (2023) explore whether ChatGPT can be a cost-effective supplement to expert feedback by acting as an automated teacher or coach.
II-B Application of ChatGPT on Other Fields
@@
Many research explore the ability of ChatGPT to understand texts in other fields than NLP. Cao et al. (2023) investigated ChatGPT’s potential cultural background by analyzing its answers to questions aimed at quantifying human cultural differences. The results of the study showed that when inspired by an American context, ChatGPT showed strong consistency with American culture, but it was less effective in adapting to other cultural contexts. George and George (2023) explore how ChatGPT can enhance e-commerce through chat and other sectors such as education, entertainment, finance, health, news, and productivity, and how ChatGPT can be used to create more personalized content for users and help businesses improve the efficiency and effectiveness of customer service. Rezayi et al. (2023) explored a new field of agricultural natural language processing by studying the effectiveness of pre-training transformer-based language models using food-related text corpora. Wang et al. (2023c) investigated ChatGPT’s ability to infer dynamic network structures from temporal text data, especially financial news. Khondaker et al. (2023) conducted a large-scale evaluation of ChatGPT on a wide range of Arabic NLP tasks. Wu et al. (2023b) evaluated ChatGPT’s ability in knowledge in the medical field on the Chinese National Medical Licensing Examination (CNMLE).
II-C Application of Large Language Models in Ancient Chinese
@@
However, there is relatively little research on ChatGPT’s understanding of ancient Chinese. The research on the understanding of ancient Chinese by language model is mainly based on BERT model. Wang et al. (2022) and Chang et al. (2023) built a pre-trained language model of SikuBERT and SikuGPT for ancient text intelligent processing tasks based on BERT deep language model framework using the verified high-quality full-text corpus of SikukuQuanshu as a training set. Some researchers built GuwenBERT based on RoBERTa model trained on a large number of ancient literature corpus. Tsinghua University has constructed BERT CCPoem, a pre-training model based on BERT specifically for ancient Chinese poetry, which is trained on the complete collection of ancient Chinese poetry CCPC full v1.0 and can be used for intelligent poetry retrieval, recommendation and sentiment analysis. Wang et al. (2023d) introduced two language models, GujiBERT and GujiGPT, which are foundational models specifically designed for intelligent information processing of ancient texts.
III Experiments
@@
III-A Data preparation
The main data comes from Github open source project NiuTrans/classic-Modern [^1] The resource can be found in website:https://github.com/NiuTrans/Classical-Modern, which is a very comprehensive ancient Chinese-modern parallel corpus, including a large number of classical ancient works. This dataset divides and displays each ancient book by chapters, and the text is stored in the txt documents under each chapter. These bilingual data are provided in sentence level, and three data formats are provided: original text, translation and bilingual.
In this study, the team mainly utilized the bilingual data of the book, Shi Shuo Xin Yu, including 36 chapters, to satisfy our needs in the experiments. And most articles in this book are about biographies of famous people in ancient China, which can be used for ancient-to-modern Chinese translation and person name recognition.
Due to the input length limitation of ChatGPT, we split the whole book into 3923 sentences and each time we feed the model a sentence. For the ancient-to-modern Chinese translation task, we include all sentences, but for the ancient Chinese name recognition task, we randomly select a total of 300 sentences because of time constraint. Also we manually label Chinese names for the 300 sentences. The data is shown in Table I as fowllows.
TABLE I: A few examples used in the experiments. The sentences are extracted from Shi Shuo Xin Yu, and Chinese names are manually labeled by the authors.
@@ III-B Accesss to ChatGPT
III-C Translation of ancient Chinese by ChatGPT
Once having access to ChatGPT in Python, we can simply design a sensible prompt to perform the ancient-to-modern Chinese translation task. We tried a variety of prompt input methods, such as “Please translate the following ancient Chinese:”, “Translate the ancient Chinese after the colon into modern Chinese:”, “Translate the ancient Chinese into modern Chinese” and so on. We found that it is necessary to include “translation” and “modern Chinese” in order to perform the ancient-to-modern Chinese translation robustly, otherwise the output of ChatGPT may have redundant and wrong information. In addition, we also found that even if we input only ancient Chinese sentences without any instruction, ChatGPT can sometimes return the correct translation results in modern Chinese. For reproduction of our experiments, the python code snippet is shown in Fig. 1.
@@ III-C Translation of ancient Chinese by ChatGPT
After many trials, we choose a straightforward “将文言文翻译成现代汉语(translate ancient Chinese into modern Chinese)” as the prompt instruction for the translation task. This prompt almost never produces error information, and no superfluous information. we extract the text content that contains only the modern Chinese translation of the ancient Chinese. Due to the input length constraints set by OpenAI, it is unfeasible to command ChatGPT to translate the entire Shi Shuo Xin Yu book directly. Consequently, we use a systematic FOR loop to sequentially translate preloaded segments of the book.
To proportionately evaluate its performance, we adjust the size of the input - the amount of text that needs to be translated. We start off with a single sentence to observe how well the program can perform a basic translation task. Then, they slowly increase the input length to three, five, and eight sentences. This gradual scaling may challenge the program and test its limits, especially in maintaining cohesiveness and continuity in the translation of longer passages.
For the ancient-to-modern translation task, we evaluate the performance of ChatGPT with BLEU (Papineni et al., 2002) and BERT-Score(Zhang et al., 2019).
III-D People Name Recognition from Ancient Chinese Literature
Besides translation, we assess the competence of ChatGPT in recognizing names referenced within the ancient Chinese text. It is noteworthy that the effectiveness of ChatGPT in executing this task is contingent upon the prompt instruction. We utilize a trial-and-error approach to ascertain a suitable prompt instruction, ”已知文言文的原文如下,找出该文言文中的人名”, which means that ”Given the following ancient Chinese text, find the people names in it”. For reproduction of our experiments, the python code snippet is shown in Fig. 2.
@@ III-D People Name Recognition from Ancient Chinese Literature
The Jieba module in Python is widely employed for word segmentation and rudimentary named entity recognition in Chinese texts. Hence, for our study on the recognition of people names in ancient Chinese text, we incorporate Jieba as a baseline method for comparison. For the people name recognition task, we evaluate the performance of ChatGPT with commonly used metrics, like precision, recall and F1 score (Jiang et al., 2016).
IV Results and Analysis
@@
IV-A Assessing the Efficacy of ChatGPT in the Translation of Ancient Chinese Texts
As mentioned in the experimental setup section III, we deploy ChatGPT to convert sentences from the ancient Chinese text, Shi Shuo Xin Yu, into modern Chinese in configurations of 1, 3, 5, and 8 sentences per query. These translations are then evaluated and contrasted using quantitative metrics such as BLEU, and BERT-Score. Table II illustrates the BLEU scores resulting from the translation of ancient to modern Chinese from ChatGPT under four different configurations: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per individual query. In this table, we conduct an evaluation of the BLEU scores across three different levels of granularity: 1-gram, 2-gram, and 4-gram. From Table II, it is observed that ChatGPT renders the most proficient translations from ancient to modern Chinese when it is engaged with three sequential ancient Chinese sentences (3-sent.), garnering 22.60% at 2-gram and 20.19‰ at 4-gram Bleu scores. This suggests that a shorter contextual input enhances the translation quality from ancient to modern Chinese. Interestingly, the quality of translation does not exhibit improvement when the context increases to eight sentences, mirroring the performance with a single sentence input. Another important finding is that the BLUE scores are still very low, reaching 0 at 4-gram BLUE in all four settings, which means that ChatGPT with GPT-3.5-Turbo still performs very poorly in ancient-to-modern Chinese translation.
TABLE II: The 1-gram, 2-gram, and 4-gram BLUE scores of translated modern Chinese texts from ChatGPT under four settings: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per query.